# IT Intrusion Detection Using Statistical Learning and Testbed Measurements

### v1

@@ Abstract
Index Terms:
automated security, intrusion detection, Hidden Markov Model, Long Short-Term Memory, SNORT, forensics
@@ I Introduction
Intrusion detection systems traditionally rely on a static set of rules for discovering possible attacks. The rules are configured by domain experts, which makes the approach flexible but expensive to maintain in the face of evolving attacks and changing IT infrastructures [1, 2].
@@
The data to train the models come from traces collected on a testbed we have built at KTH [3]. It emulates the infrastructure depicted in Fig. 1 and allows us to run various attack scenarios during which we collect measurements and alerts.
@@ II Use Case
We consider an intrusion detection use case that involves the IT infrastructure of an organization as depicted in Fig. 1 [4]. The infrastructure includes a set of servers that run application services and an IDS that logs events in real-time. The operator of this infrastructure, which we call the defender, provides services to a client population and monitors the infrastructure and the services. The clients access the services through a public gateway, which is also open to an attacker. The attacker intrudes on the infrastructure and performs an attack, for instance, retrieving information, installing malicious software, etc. During an attack, the attacker executes a sequence of attack actions including network scans, vulnerability exploits, and tool installation. The defender continuously monitors the infrastructure and analyzes the IDS statistics and measurements from network components and servers. Note that the defender does not directly observe the attacker but only indirectly learns about an attack through the monitoring data, which originates from the actions of the attacker and the clients. Using this data, the defender has the following objectives:
@@ III Problem formulation and approach
The properties of HMM have been thoroughly studied in the literature and many algorithms for HMM have been developed (see Section IV). One of them is the well-known Viterbi algorithm [5], which maps a given observation sequence to the most likely sequence of hidden states.
@@
HMM is usually used with the first-order assumption, which says that the hidden state at time $t$ only depends on the hidden state at time $t-1$ and not on earlier states. In our scenario, this means that an attack action only depends on the action of the previous time step but not on earlier time steps. This is obviously an approximation of reality, and an experimental evaluation will show whether the assumption is justified.
@@
We use the categorical version of a HMM, which assumes a finite set of possible observation symbols. This assumption is made in most of the related research that uses HMM [6, 7, 8, 9, 10, 11]. Since the number of model parameters increases linearly with the number of possible observations, we attempt to keep the number small in order to limit the sample size required to accurately estimate the model and to reduce the model complexity. In our use case, the observation vector is high-dimensional: it contains 50 observation attributes, such as host metrics and IDS alert counts. Consequently, we must find a mapping from the high-dimensional observation space onto a small set of observation symbols, which are used in the HMM. In this work, we apply feature selection and clustering techniques to accomplish this mapping (see Section VI).
@@
$N$ number of hidden states
@@ IV-A Hidden Markov Model (HMM)
A HMM is a statistical model that describes a Markov process with hidden states and associated observations. It captures the probabilistic relationship between a hidden state and an observation, as shown in Fig. 3. A HMM $\lambda$ can be defined as the triple [12]:
@@
$A$ is the transition matrix between hidden states:
@@
In case only sample sequences of observations are available, the Baum-Welch algorithm [13] allows for unsupervised training. It is an expectation-maximization algorithm that attempts to maximize $P(O|\lambda)$. The computational complexity of the Baum-Welch algorithm is $O(N^{2}TIL)$ where $I$ is the number of iterations and $L$ is the number of training sequences.
@@ IV-B Long Short-Term Memory (LSTM)
LSTM [14] has a specific recurrent neural network architecture that is capable of learning long-term dependencies. For this work, we choose a LSTM architecture with the length of the input sequence equal to the length of the output sequence. The input is a sequence of measurements and the output is a sequence of attack actions. While input values are numerical, output values are categorical.
The computational complexity of training a LSTM is $O(TH^{2}+THD)$ where $H$ is the number of cells in hidden layers [15]. We use the tensorflow.keras package [16] for training and evaluation.
@@ IV-C Random Forest Classifier (RFC)
RFC [17] is a supervised ensemble learning method that aggregates the results of multiple decision trees to determine the output. In this work, we not only use RFC for predicting attack actions but also in the data preprocessing phase to reduce the number of attributes for model training.
The computational complexity of RFC is $O(DFR\log R)$ where $F$ is the number of trees, and $R$ is the number of samples [18]. We use the sklearn.ensemble.RandomForestClassifier package [19] with default parameters.
@@ V Emulating attacks and collecting measurements
We have built a testbed at KTH which emulates the infrastructure and the configuration shown in Fig. 1. On the testbed, we emulate physical hosts, network connections, the client population, as well as the attacker and the defender. For this research, we have executed thousands of attack episodes and have obtained traces from the measurements and events observed by the defender. The length of a time step during these experiments has been 30 seconds. More information about the emulation platform is available at [3].
@@ 5 Ping Scan Network Service Login
6 DVWA SQL Injection Exploit [20] DVWA SQL Injection Exploit
@@ 17 Ping Scan Install Tools
TABLE 2: Two attack types with attack action sequences. Actions that exploit vulnerabilities in specific software products are identified by the vulnerability identifiers in the Common Vulnerabilities and Exposures (CVE) database [21].
@@ VI Data preprocessing
The traces collected from running attack scenarios on our testbed contain observation vectors with 50 attributes. Interestingly, more than half of them turn out to have constant values and are therefore redundant. We apply the correlation analysis to the remaining 14 attributes, which leads us to remove the additional three attributes. Training the RFC produces the ranking of the remaining 11 attributes. Four of them are alert counts from the IDS, which is a SNORT system [1]. The rest are statistics from servers.
@@ VII-E Discussion of the evaluation results
We applied HMM and RFC to the dataset CSE-CIC-IDS2018 [22] and obtained a prediction accuracy for the intrusion start time that is very close to the one reported for our traces. (This dataset does not allow for other predictions that we performed in our evaluation.)
@@ VIII Related work
When reviewing works that apply HMM to address problems in IT security, we found two directions. First, HMM is used to profile an actor [6, 23, 24, 25, 26], for example, an attacker or a normal user in the network environment. Based on the profile, the actor can be identified based on a sequence of observations. Second, HMM is used to identify different stages or states of an attack [7, 8, 9, 10, 11, 27, 28]. In this work, we follow both directions by constructing HMM that allows us to profile attack types and by training HMM that allows us to predict intrusion start time and attack actions.
@@
In [26], the authors argue that attack-labeled flow data sets are rare for evaluating flow-based intrusion detection techniques. Therefore, they propose a generative method for generating synthetic flow-based time series data under SSH brute-force attacks. The method is based on HMM where the hidden states correspond to the attack phases and attacker activity while the observations correspond to the number of flows, the number of packets, and the number of bytes. HMM is first trained in a supervised fashion with real network data and then the trained model is used to generate synthetic data.
@@
In [9], the authors study an intrusion detection scenario where two different attackers simultaneously intrude on the system. They develop two different methods to predict the attack type and attack stage using HMM. Both methods rely on two HMMs modeling the two attacks, and one of them introduces a special observation symbol that describes the case where an observation does not relate to the considered attack type.
@@
In [11], the authors propose a method for real-time prediction of an attack stage based on HMM. They consider three stages of a DDoS attack. They collect alerts from the SNORT system and map them onto 10 observation symbols based on analysis of CVE reports [21]. The models are trained offline using the public DARPA2000 dataset LLDDOS1.0 [29]. Their evaluation shows that they can achieve similar prediction accuracy as we report in this paper for supervised learning. For unsupervised learning, however, they achieve much lower accuracy.
@@
Some works combine HMM with neural networks [30, 31, 32]. The authors of [30] introduce the Fusion Hidden Markov Model (FHMM) for profiling attacker behavior in computer networks. A FHMM entails a stochastic ensemble of HMMs to capture the diverse behavior of attackers. The authors partition the set of training sequences into least correlated subsets. For each subset, they train a HMM and aggregate the predictions from all HMMs through a neural network. The authors motivate their approach with the need for a method robust to noise. In [31] the authors propose an attack detection and classification framework for zero-day attacks. The predictions of the trained HMM are used to train a deep neural network.
@@
In [33] the authors study the method of classifying a family of attacks using a single HMM instead of using one HMM for each attack type. They argue that different attack types often have similar characteristics and thus may share some observation symbols and hidden states. In our work, we have evaluated and compared both approaches (i.e., training a HMM for each attack type vs. training a single HMM for both attack types) and found that they lead to very similar accuracy values for attack classification.
@@
Similar to HMM, the related works that apply LSTM to intrusion detection fall into two categories. First, LSTM is used as a classifier [34, 35, 36, 37]. Second, like in this paper, LSTM is used for sequence-to-sequence learning [38, 39].
@@
The authors of [34] claim that their paper is the first successful application of LSTM to intrusion detection. They model network traffic as a time series and apply LSTM to classify the connection records into either normal traffic or one of four attack categories. Different LSTM architectures are evaluated. Like [34], the authors of [37] apply LSTM to traffic classification. They evaluate different feature reduction techniques applied before LSTM and find that PCA gives the best results.
@@
[39] applies LSTM to predict a sequence of attack stages. The observations are Snort messages which are mapped onto a vector space. The method uses an encoder-decoder architecture. One LTSM network encodes a sequence of alerts into a latent feature vector and then another LSTM network decodes this vector to a sequence of predicted attack stages.
@@
RFC has been applied to intrusion detection as a classifier [40, 41, 42, 43, 44, 45, 46]. In [42], the authors propose a two-stage classifier for network intrusion detection. The first stage uses RFC to separate the traffic into the classes “normal” and “attack.” The second stage uses a different classifier to further separate the attack traffic into attack types. In [43] RFC is used to detect mobile botnets by classifying the system calls invoked by applications.
@@ X Acknowledgements
The authors are grateful to Kim Hammar and Forough Shahab for fruitful discussions around this work and for their comments on an earlier version of this paper. This research has been supported by Digital Futures.

### v2

@@ I Introduction
Intrusion detection systems traditionally rely on a static set of rules for discovering possible attacks. The rules are configured by domain experts, which makes the approach flexible but expensive to maintain in the face of evolving attacks and changing IT infrastructures [snort, ids_survey].
@@
The data to train the models come from traces collected on a testbed we have built at KTH [hammar2023learning]. It emulates the infrastructure depicted in Fig. 1 and allows us to run various attack scenarios during which we collect measurements and alerts.
@@ II Use Case
We consider an intrusion detection use case that involves the IT infrastructure of an organization as depicted in Fig. 1 [hammar2022learning]. The infrastructure includes a set of servers that run application services and an IDS that logs events in real-time. The operator of this infrastructure, which we call the defender, provides services to a client population and monitors the infrastructure and the services. The clients access the services through a public gateway, which is also open to an attacker. The attacker intrudes on the infrastructure and performs an attack, for instance, retrieving information, installing malicious software, etc. During an attack, the attacker executes a sequence of attack actions including network scans, vulnerability exploits, and tool installation. The defender continuously monitors the infrastructure and analyzes the IDS statistics and measurements from network components and servers. Note that the defender does not directly observe the attacker but only indirectly learns about an attack through the monitoring data, which originates from the actions of the attacker and the clients. Using this data, the defender has the following objectives:
@@ III Problem formulation and approach
The properties of HMM have been thoroughly studied in the literature and many algorithms for HMM have been developed (see Section IV). One of them is the well-known Viterbi algorithm [forney1973viterbi], which maps a given observation sequence to the most likely sequence of hidden states.
@@
HMM is usually used with the first-order assumption, which says that the hidden state at time $t$ only depends on the hidden state at time t1t-1 and not on earlier states. In our scenario, this means that an attack action only depends on the action of the previous time step but not on earlier time steps. This is obviously an approximation of reality, and an experimental evaluation will show whether the assumption is justified.
@@
We use the categorical version of a HMM, which assumes a finite set of possible observation symbols. This assumption is made in most of the related research that uses HMM [srivastava2008credit, arnes2006using, zhang2014application, shawly2019architectures, ghafir2019hidden, holgado2017real]. Since the number of model parameters increases linearly with the number of possible observations, we attempt to keep the number small in order to limit the sample size required to accurately estimate the model and to reduce the model complexity. In our use case, the observation vector is high-dimensional: it contains 50 observation attributes, such as host metrics and IDS alert counts. Consequently, we must find a mapping from the high-dimensional observation space onto a small set of observation symbols, which are used in the HMM. In this work, we apply feature selection and clustering techniques to accomplish this mapping (see Section VI).
@@
NN number of hidden states
@@ IV-A Hidden Markov Model (HMM)
A HMM is a statistical model that describes a Markov process with hidden states and associated observations. It captures the probabilistic relationship between a hidden state and an observation, as shown in Fig. 3. A HMM $\lambda$ can be defined as the triple [rabiner1989tutorial]:
@@
AA is the transition matrix between hidden states:
@@
In case only sample sequences of observations are available, the Baum-Welch algorithm [rabiner1986introduction] allows for unsupervised training. It is an expectation-maximization algorithm that attempts to maximize $P(O|\lambda)$. The computational complexity of the Baum-Welch algorithm is $O(N^{2}TIL)$ where $I$ is the number of iterations and $L$ is the number of training sequences.
@@ IV-B Long Short-Term Memory (LSTM)
LSTM [hochreiter1997long] has a specific recurrent neural network architecture that is capable of learning long-term dependencies. For this work, we choose a LSTM architecture with the length of the input sequence equal to the length of the output sequence. The input is a sequence of measurements and the output is a sequence of attack actions. While input values are numerical, output values are categorical.
The computational complexity of training a LSTM is $O(TH^{2}+THD)$ where $H$ is the number of cells in hidden layers [rotman2021shuffling]. We use the tensorflow.keras package [LSTM] for training and evaluation.
@@ IV-C Random Forest Classifier (RFC)
RFC [breiman2001random] is a supervised ensemble learning method that aggregates the results of multiple decision trees to determine the output. In this work, we not only use RFC for predicting attack actions but also in the data preprocessing phase to reduce the number of attributes for model training.
The computational complexity of RFC is $O(DFR\log R)$ where $F$ is the number of trees, and $R$ is the number of samples [TBcomplexity]. We use the sklearn.ensemble.RandomForestClassifier package [RFclassifier] with default parameters.
@@ V Emulating attacks and collecting measurements
We have built a testbed at KTH which emulates the infrastructure and the configuration shown in Fig. 1. On the testbed, we emulate physical hosts, network connections, the client population, as well as the attacker and the defender. For this research, we have executed thousands of attack episodes and have obtained traces from the measurements and events observed by the defender. The length of a time step during these experiments has been 30 seconds. More information about the emulation platform is available at [hammar2023learning].
@@ 5 Ping Scan Network Service Login
6 DVWA SQL Injection Exploit [dvwa] DVWA SQL Injection Exploit
@@ 17 Ping Scan Install Tools
TABLE 2: Two attack types with attack action sequences. Actions that exploit vulnerabilities in specific software products are identified by the vulnerability identifiers in the Common Vulnerabilities and Exposures (CVE) database [cve].
@@ VI Data preprocessing
The traces collected from running attack scenarios on our testbed contain observation vectors with 50 attributes. Interestingly, more than half of them turn out to have constant values and are therefore redundant. We apply the correlation analysis to the remaining 14 attributes, which leads us to remove the additional three attributes. Training the RFC produces the ranking of the remaining 11 attributes. Four of them are alert counts from the IDS, which is a SNORT system [snort]. The rest are statistics from servers.
@@ VII-E Discussion of the evaluation results
We applied HMM and RFC to the dataset CSE-CIC-IDS2018 [sharafaldin2018toward] and obtained a prediction accuracy for the intrusion start time that is very close to the one reported for our traces. (This dataset does not allow for other predictions that we performed in our evaluation.)
@@ VIII Related work
When reviewing works that apply HMM to address problems in IT security, we found two directions. First, HMM is used to profile an actor [srivastava2008credit, annachhatre2015hidden, zhao2017software, wang2004modeling, sperotto2009hidden], for example, an attacker or a normal user in the network environment. Based on the profile, the actor can be identified based on a sequence of observations. Second, HMM is used to identify different stages or states of an attack [arnes2006using, zhang2014application, shawly2019architectures, ghafir2019hidden, holgado2017real, zan2009hidden, chen2016anomaly]. In this work, we follow both directions by constructing HMM that allows us to profile attack types and by training HMM that allows us to predict intrusion start time and attack actions.
@@
In [sperotto2009hidden], the authors argue that attack-labeled flow data sets are rare for evaluating flow-based intrusion detection techniques. Therefore, they propose a generative method for generating synthetic flow-based time series data under SSH brute-force attacks. The method is based on HMM where the hidden states correspond to the attack phases and attacker activity while the observations correspond to the number of flows, the number of packets, and the number of bytes. HMM is first trained in a supervised fashion with real network data and then the trained model is used to generate synthetic data.
@@
In [shawly2019architectures], the authors study an intrusion detection scenario where two different attackers simultaneously intrude on the system. They develop two different methods to predict the attack type and attack stage using HMM. Both methods rely on two HMMs modeling the two attacks, and one of them introduces a special observation symbol that describes the case where an observation does not relate to the considered attack type.
@@
In [holgado2017real], the authors propose a method for real-time prediction of an attack stage based on HMM. They consider three stages of a DDoS attack. They collect alerts from the SNORT system and map them onto 10 observation symbols based on analysis of CVE reports [cve]. The models are trained offline using the public DARPA2000 dataset LLDDOS1.0 [darpa]. Their evaluation shows that they can achieve similar prediction accuracy as we report in this paper for supervised learning. For unsupervised learning, however, they achieve much lower accuracy.
@@
Some works combine HMM with neural networks [deshmukh2019attacker, aoudni2022cloud, jiang2005novel]. The authors of [deshmukh2019attacker] introduce the Fusion Hidden Markov Model (FHMM) for profiling attacker behavior in computer networks. A FHMM entails a stochastic ensemble of HMMs to capture the diverse behavior of attackers. The authors partition the set of training sequences into least correlated subsets. For each subset, they train a HMM and aggregate the predictions from all HMMs through a neural network. The authors motivate their approach with the need for a method robust to noise. In [aoudni2022cloud] the authors propose an attack detection and classification framework for zero-day attacks. The predictions of the trained HMM are used to train a deep neural network.
@@
In [dass2021] the authors study the method of classifying a family of attacks using a single HMM instead of using one HMM for each attack type. They argue that different attack types often have similar characteristics and thus may share some observation symbols and hidden states. In our work, we have evaluated and compared both approaches (i.e., training a HMM for each attack type vs. training a single HMM for both attack types) and found that they lead to very similar accuracy values for attack classification.
@@
Similar to HMM, the related works that apply LSTM to intrusion detection fall into two categories. First, LSTM is used as a classifier [staudemeyer2015applying, diro2018leveraging, althubiti2018lstm, laghrissi2021intrusion]. Second, like in this paper, LSTM is used for sequence-to-sequence learning [sai2019advance, zhou2021detecting].
@@
The authors of [staudemeyer2015applying] claim that their paper is the first successful application of LSTM to intrusion detection. They model network traffic as a time series and apply LSTM to classify the connection records into either normal traffic or one of four attack categories. Different LSTM architectures are evaluated. Like [staudemeyer2015applying], the authors of [laghrissi2021intrusion] apply LSTM to traffic classification. They evaluate different feature reduction techniques applied before LSTM and find that PCA gives the best results.
@@
[zhou2021detecting] applies LSTM to predict a sequence of attack stages. The observations are Snort messages which are mapped onto a vector space. The method uses an encoder-decoder architecture. One LTSM network encodes a sequence of alerts into a latent feature vector and then another LSTM network decodes this vector to a sequence of predicted attack stages.
@@
RFC has been applied to intrusion detection as a classifier [gupta2016framework, belavagi2016performance, stefanova2017network, da2017detecting, lashkari2017characterization, mcelwee2017active, choubisa2022simple]. In [stefanova2017network], the authors propose a two-stage classifier for network intrusion detection. The first stage uses RFC to separate the traffic into the classes “normal” and “attack.” The second stage uses a different classifier to further separate the attack traffic into attack types. In [da2017detecting] RFC is used to detect mobile botnets by classifying the system calls invoked by applications.
@@ X Acknowledgements
The authors are grateful to Kim Hammar and Forough Shahab for fruitful discussions around this work and for their comments on an earlier version of this paper. This research has been supported by Digital Futures and the WASP NEST program through project $\mathrm{AIR}^{2}$.